OpenStudy (anonymous):

(8-z)/(9-z)=(1)/(z-9) what is z=

OpenStudy (anonymous):

(8-z)/(9-z)=1/(z-9) (8-z)/(9-z)= -1/(9-z) (8-z)(9-z)/(9-z)=-1 8-z=-1 9=z

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Do we need to be concerned at all that z=9 makes both sides of the equation explode?

OpenStudy (anonymous):

dividing by 0 good catch. then it is no solution.